OpenStudy (anonymous):

Can anyone tell me if I did this correctly? (6x3 - 13x + 13)(2x + 3) 6x3(2x + 3) - 13x(2x + 3) + 13(2x + 3) (6x3)(2x) + (6x3)(3) - (13x)(2x) + (13x)(3) + (13)(2x) + (13)(3) 12x4 + 18x3 - 26x2 + 36x + 26x + 36 12x4 + 18x3 - 26x2 + 62x + 36

OpenStudy (anonymous):

Sorry, the exponents did not copy through (6x^3 - 13x + 13)(2x + 3) 6x^3(2x + 3) - 13x(2x + 3) + 13(2x + 3) (6x^3)(2x) + (6x^3)(3) - (13x)(2x) + (13x)(3) + (13)(2x) + (13)(3) 12x^4 + 18x^3 - 26x^2 + 36x + 26x + 36 12x^4 + 18x^3 - 26x^2 + 62x + 36

OpenStudy (anonymous):

Minor mistake 3*13=39 not 36

OpenStudy (anonymous):

aaaaaahhh ty

OpenStudy (anonymous):

brain is toast and still got a ways to go....lol

OpenStudy (anonymous):

Also -13 * 3= is negative not positive 39

OpenStudy (anonymous):

so that would be 65x instead of 62

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Not quite, you have mistakes regarding positive and negative. 12x^4+18x^3-26x^2-39x+26x+39 =12x^4+18x^3-26x^2-7x+39

OpenStudy (anonymous):

trying again (6x^3 - 13x + 13)(2x + 3) 6x^3(2x + 3) - 13x(2x + 3) + 13(2x + 3) (6x^3)(2x) + (6x^3)(3) + (-13x)(2x) + (-13x)(3) + (13)(2x) + (13)(3) 12x^4 + 18x^3 - 26x^2 - 39x + 26x + 36 12x^4 + 18x^3 - 26x^2 - 13x + 36

OpenStudy (anonymous):

where did you get the -7?

OpenStudy (anonymous):

Your last term was multiplied incorrectly: 13*3=39 not 36

OpenStudy (anonymous):

-7x was from -39x+26x They are like terms so -39+26=-7

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

I thought it was -13

OpenStudy (anonymous):

-13, my bad :P

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

final answer: (6x^3 - 13x + 13)(2x + 3) 6x^3(2x + 3) - 13x(2x + 3) + 13(2x + 3) (6x^3)(2x) + (6x^3)(3) + (-13x)(2x) + (-13x)(3) + (13)(2x) + (13)(3) 12x^4 + 18x^3 - 26x^2 - 39x + 26x + 39 12x^4 + 18x^3 - 26x^2 - 13x + 39

OpenStudy (anonymous):

at least I was on the right track with this one.....lol

OpenStudy (anonymous):

Yes, good job

OpenStudy (anonymous):

ty